whitemonsterbunny17 (whitemonsterbunny17):

¢αη ѕσмєσηє ρℓєαѕє нєℓρ мє ƒιgυяє συт тнιѕ qυєѕтιση? ιη¢ℓυ∂є ∂ιяє¢тισηѕ ση нσω тσ ƒιgυяє ιт συт тσσ, ρℓєαѕє. тнαηкѕ! (ノ^ヮ^)ノ*:・゚✧*:・゚✧

whitemonsterbunny17 (whitemonsterbunny17):

The quantity 5.85 * 10^4 is equivalent to _____. A. 5 850 000 m B. 58 500 m C. 5 840 m D. 0.000 585 m

undeadknight26 (undeadknight26):

*facepalm* just *facepalm* miss bunnies...

ganeshie8 (ganeshie8):

too many zeroes @Luigi0210 do you love zeroes ? ;)

ganeshie8 (ganeshie8):

\[\large 10^4 = 10\times 10 \times 10 \times 10 = 10000\]

ganeshie8 (ganeshie8):

just move the decimal point to right by 4 positions

OpenStudy (anonymous):